OpenStudy (anonymous):

differentiate the function. f(x) = sin(3 ln x)

OpenStudy (anonymous):

You have to use the chain rule: \[\frac{d}{dx}(\sin(3lnx))=\frac{3\cos(3lnx)}{x}\]